hope someone could help

(-i)^4 "i" is an imaginary number
What is the correct solution?
1. i
2. -i
3. 1
4. -1

Answer by scott8148(6628)

i is the square root of -1, so i^2 is -1

(-i)^4=[(-i)^2][(-i)^2] __ squaring __ (-i)^4=(i^2)(i^2) __ squaring __ (-i)^4=(-1)(-1) __ multiplying __ (-i)^4=1